Dividing Partnership Income Dylan Howell and Demond Nickles have decided to form a partnership. They have agreed that Howell is to invest $50,000 and that Nickles is to invest $75,000. Howell is to devote full time to the business, and Nickles is to devote one-half time. The following plans for the division of income are being considered: Equal division. In the ratio of original investments. In the ratio of time devoted to the business. Interest of 10% on original investments and the remainder in the ratio of 3:2. Interest of 10% on original investments, salary allowances of $38,000 to Howell and $19,000 to Nickles, and the remainder equally. Plan (e), except that Howell is also to be allowed a bonus equal to 20% of the amount by which net income exceeds the total salary allowances. Required: For each plan, determine the division of the net income under each of the following assumptions: (1) net income of $420,000 and (2) net income of $150,000. (1) (2) $420,000 $150,000 Plan Howell Nickles Howell Nickles a. $ $ $ $ b. $ $ $ $ c. $ $ $ $ d. $ $ $ $ e. $ $ $ $ f. $ $ $ $
Answer #1

Answer Page No ③ o Nebincome of $420,000 Howell Equality 420,000 - $ 210,000 Inves&ment 60,000 50,000 + 45000 X $420000 = $16Page No ③ Investment = 75000 50000+ 75000 X 420,000 $252000 Time = 0.5 1 +0:5 x 420,000 = $140,000 Interest at 10% = 75000 xPage No ③ 50,000 - Investment = 20 X 150,000 50,000 + 75000 : = $ 60,000 Time = I x 150,000 1+0.5 = $ 100,000 Interest at 10%Page. No 4 015 Time : - X 150000 1+0.5 = $50,000 profit. - (150,000 — | 2500) Ⓡafael Remain ing = 134500 x $ 55000 Plan 0 $42